3. In captivity, the average life span of a grizzly bear is 9 years more than that of a tiger. The sum of the two average life spans is 41 years. Find the average life span of a tiger in captivity.
a. Define the variables.
b. Write the expression for the life span of a grizzly bear in terms of the life span of a tiger.
c. Write an equation for the sum of their life spans in term of one variable.
d. Solve the equation for the life span of a tiger.
4. George is 12 years older than Jeannie. The sum of their ages in 5 years will be 50. How old is Jeannie now?
a. Define the variables.
b. Write an expression for the age of George in terms of Jeannie's age.
c. Write an equation for the sum of their ages in 5 years from now.
d. Solve the equation for the age of Jeannie now.
5. Mandy begins bicycling west at 20mph. Liz leaves from the same point 1 hour later bicycling west at 30mph. When will she catch up with Mandy?
a. Define the variables.
b. Make a chart.
distance=rate x time Rate Time Distance
Mandy 20 t 20t
Liz 30 t -1 30(t-1)
c. Set the distances equal to one another (they travel the same distance) and solve the equation for t.
d. What does t represent?

Answer by stanbon(48535)

3. In captivity, the average life span of a grizzly bear is 9 years more than that of a tiger. The sum of the two average life spans is 41 years. Find the average life span of a tiger in captivity.
a. Define the variables.
Let t be tiger's life span and b be the bear's life span.
---------------
b. Write the expression for the life span of a grizzly bear in terms of the life span of a tiger. g=t+9
c. Write an equation for the sum of their life spans in term of one variable.
t+t+9 = 41
------------
d. Solve the equation for the life span of a tiger.
2t = 32
t=16
---------------
4. George is 12 years older than Jeannie. The sum of their ages in 5 years will be 50. How old is Jeannie now?
a. Define the variables.
George's age is g ; Jeannie's age if j.
-------------------------
b. Write an expression for the age of George in terms of Jeannie's age.
g = j-12
---------------
c. Write an equation for the sum of their ages in 5 years from now.
g+5 + j+5 = 50
-----------------
d. Solve the equation for the age of Jeannie now.
(j-12)+5 + j+5 = 50
2j - 2 = 50
2j = 52
j = 26 yrs old
-------------------------------

5. Mandy begins bicycling west at 20mph. Liz leaves from the same point 1 hour later bicycling west at 30mph. When will she catch up with Mandy?
a. Define the variables
Mandy's time is t; Liz's time is t-1.
b. Make a chart.
distance=rate x time Rate Time Distance
Mandy: rate = 20 mph ; time = t ; distance= 20t
Liz..: rate = 30 mph ; time = t -1 ; distance = 30(t-1)
c. Set the distances equal to one another (they travel the same distance) and solve the equation for t.
20t = 30(t-1)
20t = 30t-30
10t = 30
t = 3 hrs
============================
d. What does t represent?
Mandy's time.
=================
Cheers,
Stan H.